OpenStudy (anonymous):

how do i distributive 24p + 54q to factor

OpenStudy (anonymous):

i need the answer to 24p+54q

OpenStudy (tkhunny):

Have you considered factoring the numeric portions? In particular, 24 = 2*2*2*3 54 = 2*3*3*3 What common factors are there? That is what you need.